For a reaction with the rate constant of 0.6sec-1 and initial concentration of reactant of 0.2mol/dm3,what is the concentration of reactant after 0.5min?
1
Expert's answer
2020-08-25T12:14:30-0400

Since the rate constant is s – 1, this is a first order reaction. Consequently, the concentration of the reactant at time t can be calculated by the formula:

c = c0*e-kt, where k is the rate constant, t is the time, and c0 is the initial concentration.

So, c = 0.2 * e-0.6*30 = 3.04*10-9 mol/dm3